OpenStudy (anonymous):

The temperature T of a given mass of gas varies inversely with its volume V. The temperature of 90 cm3 of a certain gas is 30°C. What will be the temperature of the gas when it is compressed to a volume of 18 cm3? (1 point)

OpenStudy (anonymous):

6°C 150°C 120°C 75°C

OpenStudy (lgbasallote):

oops this seems to be the simpler version

OpenStudy (lgbasallote):

i did not read the whole question...lemme do the proportion equation instead

OpenStudy (lgbasallote):

inversely proportional so \[V = \frac{k}{T}\]

OpenStudy (lgbasallote):

\[90 = \frac{k}{30}\] \[k = 270\] we use this proportionality constant for the second equation

OpenStudy (lgbasallote):

\[18 = \frac{270}{T}\]

OpenStudy (lgbasallote):

\[T = \frac{270}{18}\]

OpenStudy (apoorvk):

AAAAAAAAAAAAAAAAAAAAA!!!!!! :O Actually, \[\large V_1 T_1 = V_2 T_2\] (pressure not needed - this not chemistry -_- #methinks)

OpenStudy (lgbasallote):

can you solve the rest @Kittta ??

OpenStudy (lgbasallote):

sorry for my initial mistake....when i saw volume and temperature i assumed chem

OpenStudy (anonymous):

Ii dont even understand what you guys are talking about? could you explain better

OpenStudy (lgbasallote):

volume is inversely proportional to temperature...that means \[V \alpha \frac{1}{T}\] where V is volume and T is temperature got it?

OpenStudy (lgbasallote):

but we need an equation so we change 1 into k<---this is a proportionality CONSTANT \[V = \frac{k}{T}\]

OpenStudy (lgbasallote):

got a qustion?

OpenStudy (ajprincess):

I thnk u need to convert the temperature from C to K

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

nahh this isnt chem so i guess we dont need it to be in K

OpenStudy (lgbasallote):

in the first situation it says that when Volume is 90 temperature is 30...so we plug this in \[90 = \frac{k}{30}\] if you notice k is missing so we find that out...we multiply both sdes by 30 \[k = 90 \times 30\] \[k = 270\]

OpenStudy (lgbasallote):

oops wait \[k = 2700\]

OpenStudy (anonymous):

is it 150 C?

OpenStudy (lgbasallote):

now in the second situation it says that volume is now 18 cm3 so temperature is missing because note that k is a CONSTANT so it is still the same for the second equation...therefore we have \[18 = \frac{2700}{T}\] multiply both sides by T \[18T = 2700\] now divide both sides by 18 \[T = \frac{2700}{18}\] i missed a zero in the equation i posted earlier

OpenStudy (anonymous):

150?

OpenStudy (lgbasallote):

yep :D

OpenStudy (lgbasallote):

dont forget degree celsius

OpenStudy (anonymous):

Okay thanks so much! :) Sorry I not good at this i have add :(

OpenStudy (anonymous):

*i'm